OpenStudy (anonymous):

find the equation of the linehaving the given slope and containing the given point. m=6/7,(1,-9) The equation of the line is y= ?

OpenStudy (anonymous):

Point intercept form. For point (x1, y1)\[y-y1=m(x-x1)\]where m is the slope

OpenStudy (anonymous):

solve for y

OpenStudy (abmon98):

y+9=6/7(x-1) 7y+63=6x-6 7y=6x-69

OpenStudy (abmon98):

y=6/7-9.8